a long distance trucker traveled 96 miles in one direction during a snow storm. the return trip in rainy weather was accomplished at double the speed and took 3 hours less thin. find the speed going.
Answer by mananth(16946)

d= 96
speed outward x mph
return speed = 2 x mph
time outward = 96 / x
time return = 96 / 2 x

time outward – time return = 3 hours

96/x-96/2x= 3
LCD= 2x
multiply by LCD
192-96 =2x*3
96=6x
x=16 mph